OpenStudy (anonymous):

3(x + 4) - 4(x + 3) = 7x + 16. please help me if you can

OpenStudy (anonymous):

(3x+12) - (4x + 12) = 7x+16 -x = 7x+16 -8x=16 x= -2

OpenStudy (anonymous):

hope it helps!